Find the solution:


2(dS)/(dt) - S/t = 5t^3 S^3


1
Expert's answer
2022-03-07T17:07:06-0500

The given equation can be rewritten as, "\dfrac{dS}{dt} - \dfrac{S}{2t} = \dfrac{5}{2}t^3 S^3" which is a Bernoulli's equation.


Dividing the equation by "S^3" we get "\dfrac{1}{S^3}\dfrac{dS}{dt} - \dfrac{1}{2S^2t} = \dfrac{5}{2}t^3"


Put "z = \dfrac{1}{S^2}". Then "\dfrac{dz}{dt} = -\dfrac{2}{S^3}\dfrac{dS}{dt}" and hence the equation becomes,


"\dfrac{dz}{dt} + \dfrac{2v}{t} = -5t^3" which is linear in "z".


The solution is given by,


"\begin{aligned}
ze^{\int Pdt} &= \int Q e^{\int Pdt} dt + c\\
ze^{\int \frac{2}{t}dt} &= -\int 5t^3 e^{\int \frac{2}{t} dt} dt + c\quad(P = \frac{2}{t}; Q = -5t^3)\\
zt^2 & = -\int 5t^3 \cdot t^2 dt + c\\
zt^2 & = -\frac{5t^6}{6} + c\\
z &= -\frac{5t^4}{6} + \frac{c}{t^2} = \frac{-5t^6+6c}{6t^2}\\
\frac{1}{S^2} &= \frac{-5t^6+6c}{6t^2}\qquad (\text{Replacing ~} z = \frac{1}{S^2})\\
S^{2} &= \frac{6t^2}{-5t^6+6c}\\
\therefore S& = \pm\sqrt{\frac{6t^2}{-5t^6+6c}}
\end{aligned}"